OpenStudy (anonymous):

What are the roots of the equation 2x squared + 7x-3=0

OpenStudy (anonymous):

\(\large (2x)^2+7x-3=0 \) or \(\large 2x^2+7x-3=0 \) ???

OpenStudy (anonymous):

i'm guessing u mean the second one?

OpenStudy (anonymous):

yes the second one.

OpenStudy (anonymous):

formula for finding roots of an equation : (-b + root( b square - (4*a*c)))/2a and (-b - root( b square - (4*a*c)))/2a so 2xsquare + 7x- 3=0 is of the form ax square+ bx+c=0 so a = 2; b = 7 ; c= -3 so (-7 + root 83)/4 and (-7 - root 83)/4 are the roots for the above equation

OpenStudy (anonymous):

hope it helped or need further explanation .. let me know

OpenStudy (anonymous):

Yes thank you soooo muchh!

OpenStudy (anonymous):

I have another similar question. Wanna help ?